If the cost per kilometre is $1.60, average utilisation 20 passengers per vehicle kilometre and average trip distance 10 kilometres what is the level of bus kilometres required to service this market?
1
Expert's answer
2015-06-04T00:00:45-0400
If P = $1, Qd = 250-60P, the cost per vehicle kilometer is N$1.60, average utilization au = 20 passengers per vehicle kilometer and average trip distance atd = 10 kilometers, then the level of bus kilometers required to service this market is s = Qd/au*atd = 190/20*10 = 85 thousands km.